Status Report

NASA Space Station On-Orbit Status 31 Oct 2003

By SpaceRef Editor
October 31, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 31 Oct 2003

All ISS systems continue to function nominally, except as noted previously or below.

FE Alex Kaleri prepared equipment and work area for the first session with the Russian biomedical “Pilot” experiment (MBI-15), which requires a work table, ankle restraint system and control handles for testing piloting skill.  The experiment was then conducted by both Kaleri and CDR/SO Mike Foale.  Sasha later deactivated, disassembled and stowed the gear.   [Mike’s program consists of three flight control modes (fixed, slow and fast free-flyer), each one five times, after checkout and calibration of the control handles.  Results were later reported to the ground.]

On their third day on board, both crewmembers worked out according to the full regular daily physical exercise program of 2.5 hrs each on TVIS treadmill (aerobic), RED exerciser (anaerobic) and VELO cycle with load trainer. 

The Science Officer then performed the periodic transfer of accumulated exercise data files from the workout equipment to the MEC (medical equipment computer) via PCMCIA memory card and RED log entries, for downlink on OCA comm.

Later, Mike also transferred the exercise files of the wrist-band HRM (heart rate monitor) receiver stations to the MEC for downlink, then deleted them on the HRM.  

Foale’s originally scheduled set-up of the EarthKAM payload was rescheduled for next Monday, 11/3.

Alex Kaleri terminated the bake-out cycle on the BMP micropurification system’s channel 2, moding the channel back to Purify.  After yesterday’s termination of regeneration on channel #1, this restored both filter beds to Purify mode.  [The regeneration of the air purifier filter beds is repeated every 20 days.  Each bakeout to space vacuum takes about 24 hours.]

Mike performed the regular once-a-week maintenance reboot on the operational PCS laptops.

Sasha conducted the routine technical maintenance of the SM’s SOZh environment control & life support system, incl. the toilet equipment (ASU).   Mike took care of the regular routine status checkup of autonomous Lab payloads (i.e., on PCG-STES010 verify fan running, muffler screen free of debris, clean if required).

Kaleri also conducted the weekly IMS (inventory management system) tagup with IMS specialists at MCC-M, discussing open issues concerning identification of equipment and storage locations.

CDR/SO Foale completed the periodic inspection of emergency lighting power sources (ELPS) in Node “Unity”.  Also planned for today was inspection of the ELS power supplies in Airlock “Quest” and Lab “Destiny”.

As part of regular monthly preventive maintenance of Russian Segment air ventilation systems, Kaleri worked one hour in the Funktsionalnyi-Grusovoi Blok (FGB) for a cleanup of the vent grills of the three SOTR (thermal control system) gas-liquid heat exchangers (GZhT).  Later, he also cleaned the protective mesh screens of the FGB’s ventilation fans (TsV).  The fans were powered off for the task from the ground, later turned on again.

Mike called down the “ad hoc” O2 partial pressure of the cabin air.  [O2 data for trending analyses by the ground are collected daily by the crew with the U.S. CSA-CP (compound specific analyzer-combustion products).]

The CDR also took CO2 readings with the CDMK (carbon dioxide monitoring kit).

Mike Foale deactivated the CGBA (commercial generic bioprocessing apparatus) after its cool-down period, following removal of the NASDA/GCF payload last Monday (10/27). 

The crew was given the go-ahead to disconnect the Cupola RWS DCP (robotics workstation’s display & control pane) bypass power cable at their leisure.

The two crewmembers again had a “free” hour each for station familiarization and acclimatization, as is standard daily rule for the first two weeks after starting station residency.

The momentous solar proton event terminated at 8:36pm EST last night.  The geomagnetic storm then subsided to below-threshold levels at 10:00pm EST.  [Geomagnetic storming continues at reduced levels, but not atlevels requiring continuous radiation monitoring support on the ground.  Solar event coverage was terminated, with return to normal daily shift support.]

Today’s CEO (crew earth observation) targets, limited in the current XPOP attitude by flight rule constraints on the use of the Lab nadir/science window, and including the targets of the Lewis & Clark 200-year memorial locations, were Miami (looking right for the conurbation), Wildfire smoke plume (Dynamic event:  Smoke from the three great California fires is blowing inland.  As a consequence, people with respiratory problems in Arizona have been encouraged to stay indoors.  Looking right from this pass over Arizona for oblique views of the entire region of burning and the associated smoke plumes), Monterrey, Mexico (looking left for views of Mexico’s main heavy industry center), Nairobi, Kenya (looking a touch right.  One image should be sufficient to capture the city), Mount Kilimanjaro, Kenya (nadir pass.  At present melting rates this equatorial ice cap on Kilimanjaro will disappear within two decades.  The site is small and relatively difficult to identify.  This is a long term monitoring site.  Detailed images were requested), Dar es Salaam, Tanzania (nadir pass over this ancient Indian Ocean port.  Dar es Salaam lies at the western end of the great trade network which connected Africa, the Middle East, India and Indonesia before the Portuguese expansion), and Kinshasa, Zaire (about as clear as the weather becomes.  Looking a touch left on the south side of the Congo River, at the point where the river widens to form a circular lake known as Stanley Pool).

CEO images can be viewed at the websites

See also the website “Space Station Challenge” at

SpaceRef staff editor.